OpenStudy (anonymous):

A candy box is in the shape of a regular 6 sided polygon. What is the measure of each interior angle of the candy box? A. 30 degrees B. 60 degrees C. 90 degrees D. 120 degrees ??????

OpenStudy (tanner23456):

total angle measurement for and n-gon is: 180 * (n-2) therefore we have 180 * (6-2) 180*4 total angle measurement = 720 now dived this by 6 to get the measurement of and individual angle. = 720/6 = 120